OpenStudy (boxman61):

Graph and label center and 4 points 61x^(2)+9y^(2)=144

OpenStudy (anonymous):

just divide both sides by 144.... to get it in standard form.... the center of this ellipse is 0,0. replace x=0 to get your y-intercepts replace y=0 to get your x-intercepts

OpenStudy (boxman61):

thank you that has been confusing me

OpenStudy (anonymous):

do you have the 4 points then ?

OpenStudy (boxman61):

not yet just started working on it

OpenStudy (boxman61):

so i did\[16x ^{2}+9y ^{2}= 144\]\[16x ^{2}+9y ^{2}/144 =144 /144\]\[x ^{2}/9+y ^{2}/16=1\]center of (0,0)\[\sqrt{9}=3 \]so Right and Left 3\[\sqrt{16}=4\] up and down 4 gives me point at (-3,0),(3,0),(0,4),(0,-4)

OpenStudy (anonymous):

hmm... the problem in your original post says 61... but you did the process correctly anyway...

OpenStudy (boxman61):

lol sorry it was 16

OpenStudy (anonymous):

sawright... you got it right... good work man... :)

OpenStudy (boxman61):

thank you again